The following output from MINITAB presents the results of a hypothesis test. Test of mu=49 vs. not=49, The assumed standard deviation 10.4, N 58, Mean 46.32, SE 1.412088, 95%CI (43.5523077, 49.0878923, Z -1.962526, P 0,049701, Do you reject H0 at the


1
Expert's answer
2022-06-13T17:54:06-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=49"

"H_1:\mu\not=49"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{46.32-49}{10.4/\sqrt{58}}=-1.962526"

Since it is observed that "|z|=1.962526>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-1.962526)=0.049701," and since "p=0.049701<0.05=\alpha," it is concluded that the null hypothesis is rejected.Therefore, there is enough evidence to claim that the population mean "\mu" is different than 49, at the "\alpha = 0.05" significance level.


"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{10.4}{\sqrt{58}}\approx1.365587"

The corresponding 95% confidence interval is


"CI=(\bar{x}-z_c\times SE, \bar{x}+z_c\times SE)"

"=(46.32-1.96\times 1.365587,"

"46.32+1.96\times 1.365587)"

"=(43.6434497, 48.9965503)"